OpenStudy (anonymous):

"Simple" systems of equations. I saw this one in my Diff Equ book and cannot figure it out.Normally I just do things like this with my simple understanding of lin. alg. But no clue how to do it with trig functions. If you could help explain it to me I would be very greatful. Acos(x)+Bsin(x)=0 A(-sin(x))+bcos(x)=tan(x) The answer is a=-sin(x)tan(x)=cos(x)-sec(x) and b=cos(x)tan(x)=sin(x)

OpenStudy (anonymous):

So you have the two equations: (1) Asin(x) + Bcos(x) = 0 (2) -Asin(x) + Bcos(x) = tan(x) Solve (1) for A: (3) A = -Btan(x) Substitute (3) into (2): Btan(x) + Bcos(x) = tan(x) Simplify: B(1/cos(x)) = tan(x) This gives B = cos(x)tan(x) = sin(x) as required... Back-substitute into (3) and simplify

OpenStudy (anonymous):

Wow, I guess I had a major brain fart there. Didn't see that tan(X)sin(x)+cos(x)=sec.. Thanks!

OpenStudy (anonymous):

sometimes trig functions start dancing around on the page after you've looked at them for a while... :) glad i could help.